Solve 5(3-2x)+3 less than or equal to -2(2-3x)+4. Express your answer using interval notation.
Answer by edjones(7311)

5(3-2x)+3 <= -2(2-3x)+4
15-10x+3<=-4+6x+4
-10x+18<=6x
add 10x to each side: 16x>=18
divide 16 into both sides: x>=18/16 or 9/8
X[9/8,infinity)
Ed